OpenStudy (anonymous):

Will give medal and fan The sides of the cubes are numbered 1 through 6. They are both tossed. What is the probability that they will both be a 2? A: 1/3 B: 1/6 C: 1/36

OpenStudy (anonymous):

C

OpenStudy (anonymous):

1/36

OpenStudy (anonymous):

the probability that one shows up "2" is \(\frac{1}{6}\) the same for the other the probability that both show up "2" is what you when you multiply those numbers together i.e. compute \(\frac{1}{6}\times \frac{1}{6}\)

OpenStudy (anonymous):

Since there are two cubes and you're trying to find the probability that the outcome will Be the same multiply 6 by 6 since they are both numbered 1 through 6

OpenStudy (anonymous):

ty all!